Operations Report – May 25th

Crew 195 Operations Report 25.05.18

SOL: 5

Name of person filing report: Dana Levin

Non-nominal systems: Continuing run Diesel generator at night. No service rep for Solar yet.

Notes on non-nominal systems: n/a

Generator (hours run): Overnight approximately 8 hours, on solar until 1600, now back on generator

Solar— SOC 10%

Diesel – Over 50%

Propane – 50 PSI

ATV Fuel (Ethanol free gasoline) – 1 gallon.

Water (trailer) – 485 gallons.

Water (static) – Approximately 90

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 1351243

Toilet tank emptied: No
Date last emptied: 5.22.18

ATV’s (Honda, 300, 350.1, 350.2, 350.3): all except Honda and 300

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: Approximately 1 gallon

# Hours the ATVs were Used today: 1 hours

Notes on ATVs: Honda and ATV 300 out of service.

Deimos rover used: yes
Hours: 1 hour
Beginning charge: 85%
Ending charge: 80 %
Currently charging: Yes

Curiosity rover used: Yes

Hours: 1 hour
Beginning charge: 99%
Ending charge: 85 %
Currently charging: Yes

Spirit rover used: no

Hab Car used?: No
Where was Hab Car used?:
Why was Hab Car used?:

General notes and comments:

Summary of internet: Working well

Summary of suits and radios: working well

Summary of Hab operations: No problems

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of health and safety issues: none noted

Questions, concerns and requests to Mission Support: None.

Copyright © The Mars Society. All rights reserved. | Main Site